### v1

# Geometry-aware 4D Video Generation for Robot Manipulation

Abstract
Understanding and predicting the dynamics of the physical world can enhance a robot’s ability to plan and interact effectively in complex environments. While recent video generation models have shown strong potential in modeling dynamic scenes, generating videos that are both temporally coherent and geometrically consistent across camera views remains a significant challenge. To address this, we propose a 4D video generation model that enforces multi-view 3D consistency of videos by supervising the model with cross-view pointmap alignment during training. This geometric supervision enables the model to learn a shared 3D representation of the scene, allowing it to predict future video sequences from novel viewpoints based solely on the given RGB-D observations, without requiring camera poses as inputs. Compared to existing baselines, our method produces more visually stable and spatially aligned predictions across multiple simulated and real-world robotic datasets. We further show that the predicted 4D videos can be used to recover robot end-effector trajectories using an off-the-shelf 6DoF pose tracker, supporting robust robot manipulation and generalization to novel camera viewpoints.
1 Introduction
Figure 1: Geometry-aware 4D Video Generation. Our model takes RGB-D observations from two camera views and predicts future 4D pointmaps in the coordinate frame of the reference view $v_{n}$. The blue pointmap is predicted from camera $v_{n}$, while the red pointmap shows the prediction from camera $v_{m}$ projected into the coordinate frame of $v_{n}$. RGB videos are predicted separately for each view. Together, the model enables geometry-consistent 4D video generation.
Understanding how the visual world changes with interactions is a key capability for intelligent robotic systems. For robot manipulation tasks, the robots need to anticipate how the environment changes by taking into account object motions or occlusions upon interactions over extended time horizons. Recent advancements in video generation models offer a powerful paradigm for learning such dynamic visual models: by forecasting future observations, robots can simulate possible outcomes, plan actions, and adapt to new environments.
However, it remains a challenge to generate realistic and physically plausible videos which smooth and precise robot policies can be extracted from. There are two challenges: temporal coherence, ensuring smooth, causally consistent motion over time; and 3D consistency, preserving object geometry and spatial correspondences across different viewpoints. Most existing video generation models capture one at the expense of the other. Pixel-based models [1, 2] trained on RGB videos often excel at short-term motion but lack an understanding of 3D structures, which leads to artifacts like flickering, deformation, or object disappearance. In contrast, 3D-aware approaches enforce geometric constraints but are limited to simple, static backgrounds and struggle to scale to realistic, multi-object manipulation scenarios [3, 4, 5].
In this work, we present a video generation framework that bridges this gap by unifying strong temporal modeling with robust 3D geometric consistency. Our method produces 4D videos that can be rendered into RGB-D sequences that are both coherent over time and spatially consistent across camera views. To achieve this, we introduce a geometry-consistent supervision mechanism inspired by DUSt3R [6] and adapt it for the video generation task. Specifically, the model is trained to predict a pair of 3D pointmap sequences: one for a reference view and one for a second view projected into the reference view camera coordinate frame. By minimizing the difference between reference and projected 3D points over time, the model learns a shared scene representation across views. This enables robust generalization to novel viewpoints at inference, which is particularly useful in robotic applications where even small camera view shifts can push visuomotor policies out of distribution and lead to failures.
Pretrained video diffusion models provide strong visual and motion priors learned from large-scale video datasets. To enhance temporal coherence, we initialize our model with pretrained weights and extend it to jointly generate future RGB frames and pointmaps. The RGB frames are trained using the original video generation loss, while the pointmaps are supervised using the proposed geometry-consistent loss. This combination enables the model to leverage the temporal priors of pretrained models while enforcing spatial and cross-view consistency through pointmap alignment, resulting in spatio-temporally consistent RGB-D video generation. We evaluate the 4D video generation quality on both simulated and real-world tasks, and our approach outperforms baselines in both video quality and cross-view consistency.
We further demonstrate that the predicted multi-view RGB-D videos can be directly used to extract robot end effector trajectories using an off-the-shelf 6DoF pose tracker, such as FoundationPose [7]. We evaluate this approach on three simulated robot manipulation tasks where the camera views are unseen during training, achieving good success rates on all tasks and outperforms the baseline method. Additionally, our generation model with geometry-consistent supervision allows any novel view to be projected into the known reference frame without additional camera calibration, enabling flexible camera placement and simplifying robot deployment.
In summary, we propose a 4D video generation framework that achieves both 3D geometric consistency and temporal coherence. To achieve this, we first introduce a geometry-consistent supervision mechanism that enforces cross-view alignment of generated videos over time. Second, we develop a benchmark for video generation in robotic manipulation, comprising both simulation and real-world tasks. Each task is recorded from diverse camera viewpoints, enabling comprehensive evaluation of 4D generation quality and generalization to unseen views. Finally, we demonstrate that the generated 4D videos can be directly used to extract robot trajectories using an off-the-shelf 6DoF pose tracker, enabling reliable manipulation under novel viewpoints without additional camera calibration.
2 Related Work
Video Generation has been a long-standing task in computer vision. Early works use recurrent networks [8, 9] or generative adversarial networks [10] to learn temporal dynamics from video data. With the recent success of image diffusion models, many works have extended diffusion models for the video prediction task by adding additional temporal layers [1, 2] and using latent diffusion techniques [11, 12, 13, 14]. In addition to RGB videos, video diffusion models have also been shown to generate other modalities such as high-fidelity depth videos [15, 16, 17] and pointmaps [18]. In this work, we adapt a latent video diffusion model (SVD) to jointly learn from RGB and depth videos with improved temporal and spatial consistency.
Multi-View and 4D Video Generation. Recent advances in camera-conditioned video generation improve spatial consistency by associating camera poses with multi-view videos [19, 20, 21, 22]. However, 4D video generation requires joint reasoning about both object geometry across views and motion over time. Prior 4D generation works separately optimize for temporal consistency with a video model and spatial consistency with a novel view synthesis model [23, 24, 25]. In this work, we present a 4D video generation framework unifies the two objectives. In particular, our method introduces a geometry-consistent supervision mechanism through cross-view pointmap alignment, inspired by DUSt3R [6], on top of the standard video diffusion process. This approach enforces spatial consistency across camera views while ensuring temporal coherence. Our goal aligns with recent 4D generation works that jointly optimize spatial and temporal consistency [3, 4], but while they primarily focus on single-object videos with white backgrounds, we target multi-object, dynamic robot manipulation scenes.
Generative Models for Robot Planning. With the recent success of video generation models and their strong generalization capabilities across diverse visual scenes, many works have explored their potential as dynamics models for robotic tasks. Specifially, robot actions can be extracted from predicted future frames using a learned inverse dynamics model [26, 27, 28], a behavior cloning policy conditioned on generated outputs [29, 30, 31], or an RGB-based pose tracking model [32]. To more tightly couple future state prediction with action inference, recent methods have proposed unified models that simultaneously predict both future video frames and robot actions [33, 34]. Yet it remains a challenge to generate spatially consistent predictions across views and capture accurate 3D geometry needed for precise robot manipulation. To bridge this gap, we propose a model that learns spatial correspondences across camera views, enabling better RGBD video generation quality on novel views and improved end effector pose tracking accuracy for better robot task performance.
3 Method
Figure 2: 4D Video Generation for Robot Manipulation. Our model takes RGB-D observations from two camera views, and predicts future pointmaps and RGB videos. To ensure cross-view consistency, we apply cross-attention in the U-Net decoders for pointmap prediction. The resulting 4D video can be used to extract the 6DoF pose of the robot end-effector using pose tracking methods, enabling downstream manipulation tasks.
In this section, we present a 4D video generation algorithm that unifies temporal coherence with 3D geometric consistency, enabling spatio-temporally consistent predictions over time and across varying viewpoints.
Problem Statement. Traditional video generation models predict future RGB frames $\{{\mathbf{O}_{t+1},\cdots,\mathbf{O}_{t+h}}\}$ from past observations $\{{\mathbf{O}_{t-h^{\prime}+1},\cdots,\mathbf{O}_{t}}\}$ captured from a single view, where $h$ and $h^{\prime}$ are the history and future horizon respectively. However, single-view prediction lacks geometric grounding and often results in temporally plausible but spatially inconsistent outputs. To address this, we introduce a stereo-based video generation network with additional 3D supervision to enforce consistent scene geometry across time and views. For a pair of video frames $\mathbf{O}_{t}^{m}$ and $\mathbf{O}_{t}^{n}$ taken at time $t$ from camera views $v_{m}$ and $v_{n}$, each pixel coordinate $(i,j)\in\{1,\cdots,W\}\times\{1,\cdots,H\}$ corresponds to a 3D point computed from its depth. This results in pointmaps $\mathbf{X}_{t}^{m}$ and $\mathbf{X}_{t}^{n}$, where $\mathbf{X}_{t}^{n}\in\mathbb{R}^{W\times H\times 3}$ represents the per-pixel 3D coordinates for view $v_{n}$, and similarly for $\mathbf{X}_{t}^{m}$. Here, $W$ and $H$ denote the image width and height, respectively.
We first describe our video diffusion backbone for generating RGB videos (§ 3.1), next we introduce our geometry-consistent supervision mechanism on pointmaps across views to enforce spatial alignment (§ 3.2). We integrate temporal dynamics with 3D geometric consistency through joint optimization (§ 3.3). Finally, we demonstrate how 4D video predictions can be used to recover robot end effector poses $\mathbf{T}_{t}\in\text{SE}(3)$ at each time step using off-the-shelf trackers (§ 3.4).
3.1 Diffusion-Based Video Generation
We adopt the Stable Video Diffusion [11] framework which has demonstrated strong performance in generating short, temporally coherent video sequences. It first projects historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ into a latent space using a pretrained Variational Autoencoder (VAE) [35] encoder. A diffusion model, implemented as a U-Net with an encoder-decoder structure, then predicts future latent representations $\{\mathbf{z}_{t+1},\cdots,\mathbf{z}_{t+h}\}$, which are decoded back into RGB frames $\{\mathbf{O}_{t+1},\cdots,\mathbf{O}_{t+h}\}$ using VAE decoder. The diffusion model $f_{\theta}$ is trained using an alternative of the standard DDPM [36] method, which directly predicts the original clean data from the noisy input at each diffusion step. The training objective for predicting a future latent $z_{t^{\prime}}$ at timestep $t^{\prime}$ is to minimize:
$\begin{array}[]{c}\mathcal{L}_{\text{diff}}(t^{\prime})=\ \mathbb{E}_{\epsilon_{t^{\prime}},\mathbf{z}_{t^{\prime}}(0),k}\left[\left\|\mathbf{z}_{t^{\prime}}(0)-f_{\theta}(\mathbf{z}_{t^{\prime}}(k),k)\right\|^{2}\right]\\ \text{where }\mathbf{z}_{t^{\prime}}(k)=\sqrt{\alpha_{k}}\,\mathbf{z}_{t^{\prime}}(0)+\sqrt{1-\alpha_{k}}\,\epsilon_{t^{\prime}},\quad\epsilon_{t^{\prime}}\sim\mathcal{N}(0,I)\end{array}$ (1)
Here $\epsilon_{t^{\prime}}$ denotes Gaussian noise, $\mathbf{z}_{t^{\prime}}(0)$ denotes the un-noised latent, and $\mathbf{z}_{t^{\prime}}(k)$ is the noised latent at diffusion step $k$. During inference, videos are generated by progressively denoising random Gaussian noise using the trained diffusion model.
3.2 Geometry-Consistent Supervision
To enforce 3D consistency across views, we adopt the cross-view pointmap supervision strategy from DUSt3R [6], adapted to the video generation setting. As shown in Figure 2, given the history pointmaps $\{\mathbf{X}_{t-h+1}^{n},\cdots,\mathbf{X}_{t}^{n}\}$ from camera view $v_{n}$, we first encode them using a Pointmap VAE, which is initialized from the pretrained RGB VAE from SVD [11] and fine-tuned on pointmap data. This produces the latent representation $\{\mathbf{z}_{t+1}^{n},\cdots,\mathbf{z}_{t+h}^{n}\}$. We then apply the same latent diffusion method used for RGB video prediction to forecast future pointmaps in the latent space. The predicted latents are subsequently decoded by the Pointmap VAE decoder to obtain future pointmaps $\{\mathbf{X}_{t+1}^{n},\ldots,\mathbf{X}_{t+h}^{n}\}$.
In parallel, the model also predicts future pointmaps from a second camera view $v_{m}$, but instead of generating them in their native frame, it expresses them in the coordinate frame of view $v_{n}$. This results in a sequence of projected pointmaps $\{\mathbf{X}_{t+1}^{m\rightarrow n},\ldots,\mathbf{X}_{t+h}^{m\rightarrow n}\}$. Each of these predictions are encoded into latent representations that are aligned with view $v_{n}$, enabling supervision through cross-view consistency.
During training, we supervise the model $f_{\theta}$ at each future time step $t^{\prime}$ using diffusion losses applied to both the native view $v_{n}$ and the projected view $v_{m}\rightarrow v_{n}$:
$\mathcal{L}_{\text{3D-diff}}(t^{\prime})=$ $\mathbb{E}_{\epsilon_{t^{\prime}}^{n},\,\mathbf{z}_{t^{\prime}}^{n}(0),\,k}\left[\left\|\mathbf{z}_{t^{\prime}}^{n}(0)-f_{\theta}\big{(}\mathbf{z}_{t^{\prime}}^{n}(k),\,k,\,c^{n}\big{)}\right\|^{2}\right]$ (2)
$+\mathbb{E}_{\epsilon_{t^{\prime}}^{m},\,\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(0),\,k}\left[\left\|\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(0)-f_{\theta}\big{(}\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(k),\,k,\,c^{m}\big{)}\right\|^{2}\right]$
where $\mathbf{z}_{t^{\prime}}^{n}(k)=\sqrt{\alpha_{k}}\,\mathbf{z}_{t^{\prime}}^{n}(0)+\sqrt{1-\alpha_{k}}\,\epsilon_{t^{\prime}}^{n},$
$\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(k)=\sqrt{\alpha_{k}}\,\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(0)+\sqrt{1-\alpha_{k}}\,\epsilon_{t^{\prime}}^{m},\quad\epsilon_{t^{\prime}}^{n},\epsilon_{t^{\prime}}^{m}\sim\mathcal{N}(0,I)$
where $\mathbf{z}_{t^{\prime}}^{n}(k)$ denotes the noised latent of the pointmap from view $v_{n}$ at diffusion step $k$, and $\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(k)$ is the noised latent for the pointmap from view $v_{m}$ projected into the coordinate frame of view $v_{n}$. Similar to Equation 1, $\epsilon_{t^{\prime}}^{n}$ and $\epsilon_{t^{\prime}}^{m}$ are Gaussian noise added to the pointmap latents from view $v_{n}$ and $v_{m}$, respectively. The conditioning latents $c^{n}$ and $c^{m}$ are derived from the historical pointmaps of their respective views. See Appendix for model architecture details.
While camera poses are required during training to define the projection from view $v_{m}$ to the coordinate frame of $v_{n}$, a key advantage emerges at inference. Given an observation from a novel view $v_{m}$, the model can directly predict pointmaps in the coordinate frame of $v_{n}$, eliminating the need for camera poses as inputs during testing.
Multi-View Cross-Attention for 3D consistency. Unlike RGB video prediction, where each view predicts future frames independently in its own coordinate system and a single shared U-Net diffusion model can be used across views, pointmap prediction requires enforcing 3D alignment across views. The native view $v_{n}$ predicts pointmaps in its own frame, while the second view $v_{m}$ predicts pointmaps projected into the coordinate frame of $v_{n}$. To reflect this asymmetry, we use two separate decoders in the U-Net diffusion model (with identical architecture but independent weights) and introduce cross-attention layers between the decoders to enable information transfer. Specifically, the intermediate features from the decoder of view $v_{n}$ are passed to the corresponding stage in the decoder of view $v_{m}$ through cross-attention. This allows the decoder of $v_{m}$ to attend to and incorporate geometric cues from $v_{n}$, facilitating accurate pointmap prediction in the reference coordinate frame. This mechanism enhances cross-view geometric consistency, particularly under viewpoint variations.
3.3 Joint Temporal and 3D Consistency Optimization
The pretrained video diffusion model provides strong temporal priors for predicting scene dynamics, while the 3D pointmap supervision enforces 3D geometric consistency across views. We leverage the pretrained video model and optimize it with both the RGB-based video diffusion loss and the pointmap-based 3D consistency loss. The full training objective is defined as the sum of losses across all predicted time steps $t^{\prime}\in{t+1,\ldots,t+h}$ and both camera views, $v_{n}$ and $v_{m}$.
$\mathcal{L}=\sum_{t^{\prime}=t+1}^{t+h}\left[\underbrace{\mathcal{L}_{\text{diff}}^{n}(t^{\prime})+\mathcal{L}_{\text{diff}}^{m}(t^{\prime})}_{\text{RGB loss}}+\lambda\cdot\underbrace{\mathcal{L}_{\text{3D-diff}}(t^{\prime})}_{\text{pointmap loss}}\right],$ (3)
where $\lambda$ balances the contribution of the geometric supervision. We set $\lambda=1$ in our experiments. This joint objective encourages both temporal coherence and cross-view 3D consistency.
3.4 Robot Pose Estimation from 4D videos
We leverage the predicted 4D video to extract robot trajectories using an off-the-shelf 6DoF pose tracking model, FoundationPose [7]. The model takes as input RGB-D frames from a single view, a binary mask of the target object in the initial frame (generated using SAM2 [37]), camera intrinsics, and the gripper CAD model. At each pose estimation timestep, the model outputs the estimated pose of the CAD model, $\mathbf{T}_{t}\in\text{SE}(3)$, along with a confidence score for the prediction, and tracks the object for subsequent frames.
Pose estimation is run independently for both views, and the result with the highest confidence score is selected. Since all outputs are aligned to the coordinate frame of the first camera view, only the first camera’s extrinsic must be calibrated. This eliminates the need for calibrating additional cameras at test time and enables flexible multi-view configurations.
To infer the gripper open/close state, we segment the left and right gripper fingers and project their pixels into 3D space based on the predicted RGB-D sequences. The distance between the centroids of the two finger point clouds is measured: if it falls below a threshold $\delta$, the gripper is considered closed; otherwise, it is considered open. The recovered trajectories are directly used to control the robot to execute downstream tasks.
4 Experiments
Figure 3: Robot Manipulation Tasks in Simulation.
4.1 Tasks
We evaluate the 4D video generation results on three simulated tasks and one real-world task. The tasks are drawn from the LBM environment [38], a physics-based simulator built on Drake [39] that provides realistic rendering and demonstrations for both single-arm and bimanual robot manipulation.
Simulation Task: The simulation tasks are illustrated in Figure 3: StoreCerealBoxUnderShelf, PutSpatulaOnTable, and PlaceAppleFromBowlIntoBin. In StoreCerealBoxUnderShelf, a single robot arm picks up a cereal box from the top of a shelf and inserts it into the shelf below. Occlusions occur during insertion, especially from certain camera viewpoints, making multi-view predictions essential. Additionally, the pick-and-insert action requires spatial understanding and precision. In PutSpatulaOnTable, the robot arm retrieves a spatula from a utensil crock and places it on the left side of the table. This task requires precise manipulation to successfully grasp the narrow object. In PlaceAppleFromBowlIntoBin, one arm picks up an apple from a bowl on the left side of the table and places it on a shelf; a second arm then picks up the apple and deposits it into a bin on the right side. This is a long-horizon, bimanual task that tests the model’s ability to predict both temporally and spatially consistent trajectories.
The dataset consists of 25 demonstrations for the StoreCerealBoxUnderShelf task, 22 demonstrations for the PutSpatulaOnTable task, and 20 demonstrations for the PlaceAppleFromBowlIntoBin task, with different initial object configurations. Each demonstration includes RGB-D observations from 16 different camera poses. We randomly sample 12 views for training and reserve the remaining 4 for testing. All camera poses are sampled from the upper hemisphere positioned above the workstation. More details and visualization can be found in Section A.2.
Real-world Task: The real-world task mirrors the PutSpatulaOnTable task from simulation. We collected 5 demonstration videos using two FRAMOS D415e cameras positioned at different viewpoints. Each camera records synchronized RGB-D observations of the manipulation sequence. This setup allows us to evaluate the model’s ability to generate real-world robot videos.
Method Cross-view Consist. RGB Depth
mIoU ($\uparrow$) FVD-$n$ ($\downarrow$) FVD-$m$ ($\downarrow$) AbsRel-$n$ (↓) AbsRel-$m$ ($\downarrow$) $\delta_{1}$-$n$ ($\uparrow$) $\delta_{1}$-$m$ ($\uparrow$)
Task 1: StoreCerealBoxUnderShelf
OURS 0.70 411.20 561.43 0.06 0.11 0.95 0.92
OURS w/o MV attn 0.41 497.43 607.73 0.15 0.31 0.75 0.66
4D Gaussian [40] 0.39 1208.00 1094.98 0.20 0.31 0.74 0.63
SVD [11] 977.06 743.25
SVD w/ MV attn 941.73 653.44
Task 2: PutSpatulaOnTable
OURS 0.69 377.68 257.70 0.03 0.07 0.98 0.97
OURS w/o MV attn 0.44 451.54 302.29 0.10 0.33 0.89 0.41
4D Gaussian [40] 0.46 1241.13 815.77 0.33 0.30 0.43 0.37
SVD [11] 370.92 417.56
SVD w/ MV attn 536.02 445.68
Task 3: PlaceAppleFromBowlIntoBin
OURS 0.64 490.88 366.98 0.06 0.07 0.95 0.96
OURS w/o MV attn 0.26 597.05 573.73 0.14 0.49 0.76 0.30
4D Gaussian [40] 0.44 1396.10 1191.40 0.18 0.16 0.80 0.81
SVD [11] 659.52 628.01
SVD w/ MV attn 812.94 766.52
Task 4: Real World PutSpatulaOnTable
OURS 0.58 232.31 239.06 0.06 0.08 0.97 0.95
OURS w/o MV attn 0.34 266.94 254.34 0.11 0.11 0.89 0.89
4D Gaussian [40] 0.00 2002.48 2708.86 0.30 1.75 0.82 0.26
SVD [11] 293.07 319.06
SVD w/ MV attn 288.12 257.90
Table 1: Multi-view 4D Video Generation Results under Novel Camera Views. We compare our method with baselines in terms of cross-view consistency, RGB video generation quality, and depth generation quality. Our method consistently enables high-quality video and depth generation while maintaining strong cross-view consistency on both simulated and real-world datasets.
Figure 4: Qualitative Results and Comparisons under Novel Camera Views. Our method generates geometrically consistent 4D videos across camera views. In contrast, baseline results often exhibit significant cross-view inconsistencies or contain noticeable artifacts in the RGB or depth predictions.
4.2 4D Video Generation Results
Evaluation Metrics. We evaluate the proposed method and baselines across three key aspects: RGB video generation quality, depth generation quality, and cross-view 3D point consistency.
Video prediction: To evaluate RGB video generation quality, we compute the commonly used Fréchet Video Distance [41] (FVD) between the generated video and ground-truth video. FVD-$n$ evaluates the prediction from the reference view $v_{n}$, and FVD-$m$ evaluates the predcition from view $v_{m}$.
Depth prediction: To evaluate the quality of the generated depth, we extract the z-axis values from the predicted pointmaps and compare them with the ground truth depth images. We use two standard depth evaluation metrics: absolute relative error ($\text{AbsRel}=|y-\hat{y}|/y$) and threshold accuracy ($\delta_{1}=\max(\hat{y}/y,y/\hat{y})<1.25$), where $y$ is the ground truth depth and $\hat{y}$ is the predicted depth.
Cross-View 3D Consistency: To evaluate the 3D consistency of the generated pointmaps across views, we compute the mean Intersection-over-Union (mIoU) on object masks. We use SAM2 [37] to track the robot gripper and obtain binary masks from the generated videos for both views. For each frame, we lift the gripper mask in view $v_{n}$ to 3D space and then re-project it to view $v_{m}$. We then compute the IoU between the bounding boxes of the projected and original gripper masks. The mIoU is averaged over all time steps, and higher values indicating stronger 3D alignment across views.
Baselines. We compare our method with prior 4D generation approaches and variants of our model to evaluate generation quality and multi-view consistency. All models are trained on the same multi-view RGB-D video dataset and tested on novel viewpoints not observed during training.
  • OURS w/o MV attn: We remove the multi-view cross-attention mechanism in the U-Net decoder; each view is instead assigned a separate decoder with no information sharing between them.
  • 4D Gaussian [40]: A baseline method that predicts one single-view RGB video using a finetuned SVD model [11] on our dataset, and then use a 4D Gaussian method, Shape of Motion [40], to reconstruct a dynamic 4D scene from the video.
  • SVD [11]: Stable Video Diffusion is a state-of-the-art video generation model. We finetune it on our dataset to predict stereo RGB video sequences.
  • SVD w/ MV attn: We finetune SVD on our dataset to predict stereo RGB video sequences, with additional multi-view cross-attention layers added in the U-Net decoder; similar to our method, each view is assigned a separate decoder.
Results on Simulation Tasks. Results are reported in Table 1, and Figure 4 visualizes the predictions produced by different methods. All results are evaluated on novel views not seen during training. Our method OURS consistently achieves the best or highly competitive results across all tasks. For multi-view RGB video generation, it produces lower FVD scores than all baselines in most cases, indicating high temporal coherence and visual fidelity. Additionally, our method achieves the best depth prediction and cross-view consistency scores, demonstrating that the geometry-consistent supervision effectively guides the model to generate spatially aligned and geometrically coherent videos.
We show that multi-view cross-attention is a crucial design choice for helping the model learn 3D geometric correspondences across camera views. The variant without multi-view cross-attention OURS w/o MV attn exhibits significantly lower 3D consistency, as measured by the mIoU metric. The video and depth generation quality also degrades in both views. In particular, the model performs poorly on the AbsRel-$m$ and $\delta_{1}$-$m$ metrics, indicating that it fails to learn the transformation needed to generate pointmaps in view $v_{n}$’s camera coordinate frame from view $v_{m}$ without the support of cross-attention layers. In Figure 4, the gripper pose of OURS w/o MV attn is inconsistent across the two RGB views, and the projected gripper mask significantly misaligns with the actual gripper mask, as shown in the last column.
4D Gaussian performs worse in video and depth generation quality, as well as cross-view consistency, as shown in Table 1. In Figure 4, the generated RGB frames appear blurry from unseen viewpoints, and the predicted depth is inaccurate. This is because the model is optimized on one single generated RGB video with fixed camera view, making it difficult to synthesize novel views.
For the other baselines, SVD and SVD w/ MV attn, depth is not predicted and no geometric supervision is applied. While multi-view cross-attention is added to SVD w/ MV attn to enable information transfer between camera views, it operates over RGB features rather than pointmaps—which naturally encode 3D structure, as used in our method. As a result, the generated videos lack 3D geometric consistency across views. As shown in Figure 4, SVD w/ MV attn produces a noticeable gripper mismatch between the two views and the RGB generation quality is significantly worse than our method.
Results on Real-world Task. We also evaluate our method on the real-world PutSpatulaOnTable task, as shown in Table 1. We use the checkpoint trained on the same task in simulation and finetune it on 5 real-world demonstration videos. In Figure 5, we show qualitative RGB-D prediction results. The model is able to generate high-fidelity future RGB-D sequences, accurately capturing both visual appearance and depth over time comparing to the baselines.
Figure 5: Real World 4D Video Generation Results on PutSpatulaOnTable. Our model predicts high-fidelity RGB-D sequences that capture the robot gripper motions. In this particular sequence, the model correctly predicts how the robot reaches the spatula, grasps it, and lifts it up from the utensil crock.
4.3 Robot Policy Results
Method Task 1 Task 2 Task 3 Avg
Dreamitate [32] 0.10 0.17 0.00 0.09
Diffusion Policy [42] 0.10 0.27 0.00 0.12
OURS 0.73 0.67 0.53 0.64
Table 2: Task Success Rate for Manipulation Tasks.
We evaluate the robot policy’s accuracy and generalization to novel camera views across three simulated tasks. Each task is tested 30 times with varying object positions and camera viewpoints. The success rate of task completion is reported in Table 2.
The video generation model takes RGB-D observations from two novel camera views as input and predicts future observations. The generated 4D video is then passed to the pose tracking model to extract 6DoF gripper poses for both robot arms. Gripper openness is inferred using the method described in Section 3.4. The robot executes actions in an open-loop manner: after each execution, the next inference is performed using the updated RGB-D observations. Each inference takes approximately 30 seconds to generate 10 future steps on 1 NVIDIA GeForce RTX 4090 GPU.
Baselines. We train baselines on the same dataset described in Section 4.1 as our method, and test on novel camera views during deployment.
Dreamitate [32]: a state-of-the-art video generation method for visuomotor policy. Dreamitate uses a pretrained SVD model and finetunes it on robotic task videos to generate stereo RGB video predictions. Since it does not predict depth, it employs MegaPose [43] to extract the 6DoF pose of end effectors from the generated videos.
Diffusion Policy [42]: a UNet diffusion model that predicts future robot end-effector trajectories, conditioned on history RGB image observations from two camera views. We randomly sample two camera views in the training dataset and encode their corresponding RGB observations using a CLIP-pretrained ViT model [44], following the same encoder setup as in [42]. The image features are then concatenated and fed as a condition to the U-Net diffusion model to predict future actions. The diffusion model outputs robot end effector trajectories in the next 16 steps. The model is evaluated on two unseen camera views.
As shown in Table 2, Dreamitate consistently underperforms across all tasks. The lack of depth prediction and geometric consistency supervision results in lower-quality, view-inconsistent video outputs, which degrade the accuracy of the extracted poses. Consequently, the downstream robot policy suffers from a high failure rate. In contrast, our method—which jointly predicts RGB-D sequences and enforces 3D consistency—achieves over 55% higher task success rate on average.
In addition, Diffusion Policy struggles to generalize to unseen viewpoints, even though it is trained on demonstrations from multiple views. This is because the model does not explicitly model geometric correspondences across views and it’s challenging for the model to learn view-invariant actions by simply conditioning on features extracted from multi-view images.
5 Limitation
First, our method requires an RGB-D video dataset with varying camera viewpoints for training. While such datasets are easy to generate in simulation, collecting them in the real world is challenging due to hardware constraints and camera calibration requirements. Additionally, most of our experiments assume clean depth images; however, obtaining high-quality depth in real-world settings is often difficult. Recent advances in bridging the sim-to-real gap for depth data, such as FoundationStereo [45], show promising results and could be leveraged to support real-world data generation in future work. Second, the inference speed of the current video generation model is relatively slow. Recent flow matching [46, 47] or autoregressive transformers [48, 49, 50, 51] have demonstrated faster inference speed, which could lead to more reactive robot policies in future works.
6 Conclusion
We present a 4D video generation model that produces spatio-temporally consistent RGB-D sequences. Our method introduces geometric-consistent supervision during training by projecting pointmaps from one camera view into another to enforce cross-view consistency. By learning a shared geometric space, the model can generate future RGB-D videos from novel viewpoints without requiring camera poses at inference time. We demonstrate improved video generation quality and 3D consistency compared to baseline methods. Additionally, the generated 4D videos can be directly used to extract robot actions using an off-the-shelf 6DoF pose tracking model, enabling effective execution of robot manipulation tasks.
Acknowledgments
The authors would love to thank Basile Van Hoorick and Kyle Sargent for sharing their codebase for SVD fine-tuning; Ruoshi Liu and Yinghao Xu for braintorming and technical discussions. In addition, we would like to thank all REAL lab members: Yifan Hou, Hojung Choi, Mengda Xu, Huy Ha, Mandi Zhao, Xiaomeng Xu, Austin Patel, Yihuai Gao, Chuer Pan, Haochen Shi, Haoyu Xiong, Maximilian Du, Zhanyi Sun, et al. for fruitful research discussions, feedback on paper writing, and emotional support. We would like to thank interns and collaborators at Toyota Research Institute: Alan Zhao, Chuning Zhu, Zhutian Yang, Naveen Kuppuswamy, Patrick Tree Miller, Blake Wulfe, and more, for project discussions and technical support. This work was supported in part by the Toyota Research Institute, NSF Award #2143601, #2037101, and #2132519, Sloan Fellowship. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
Appendix A Technical Appendices and Supplementary Material
In Section A.1, we provide more details of our 4D generation model architecture. In Section A.2, we describe the camera sampling method used to generate the multi-view RGBD video dataset for training the 4D generation model. In Section A.3, we provide details such as compute resources requirements and hyperparameter choices for model training and inference. In Section A.4, we provide more quantitative and qualitative results of our method and additional baselines. In Section A.5, we discuss the broader impact of our work.
A.1 Model Details
In §3.1 and §3.2, we discussed that the model takes in historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ and historical pointmaps $\{\mathbf{X}_{t-h+1},\cdots,\mathbf{X}_{t}\}$ from both the native view $v_{n}$ and the second view $v_{m}$. In practice, we use the latest observation, and repeat it $h=10$ times to match the number of frames that needs to be predicted, following the implementation in SVD [11].
Each pair of RGB image and pointmap condition ${\mathbf{O}^{v}_{t},\mathbf{X}^{v}_{t}}$, where $v\in\{n,m\}$, is independently encoded using separate VAE encoders for images and pointmaps, as detailed in §3. The image VAE encodes each RGB frame into a latent feature of shape $h\times c\times w^{\prime}\times h^{\prime}$, where $h=10$ is the temporal horizon, $c{=}4$ is the latent channel size, $w^{\prime}{=}32$ and $h^{\prime}{=}40$ are spatial dimensions of the latent feature maps. Similarly, the pointmap VAE encodes pointmap into shape $h\times c\times w^{\prime}\times h^{\prime}$. These encoded image and pointmap features are then concatenated along the channel axis with the corresponding noisy latents of future images and pointmaps, yielding a combined input tensor of shape $h\times 4c\times w^{\prime}\times h^{\prime}$ ($h\times 16\times 32\times 40$), which is fed into the U-Net diffusion model.
Figure 6: Multi-View Cross-Attention. We insert a cross attention layer after each decoder block in the U-Net diffusion model for view $v_{m}$. By cross-attending to features in the native view $v_{n}$, the cross-attention layers allow information sharing between view branches.
To allow information sharing between the two diffusion branches as shown in Figure 2, we add one cross-attention layer after each decoder block in the U-Net diffusion model for the branch corresponding to view $v_{m}$. This results in 12 added cross attention layers. As illustrated in Figure 6, the query to the cross-attention layer are feature map tokens (feature at each pixel in the feature map) output by the decoder block in view $v_{m}$, where $c^{\prime}$ is the feature dimension, $h^{\prime}$ and $w^{\prime}$ are spatial dimensions of the feature map; the key and value are feature map tokens output by the corresponding decoder block in the native view $v_{n}$’s branch. The updated features are passed to the next decoder block in view $v_{m}$. The cross-attention layers capture spatial correspondences between the views through our geometric-consistent supervision mechanism.
We use pre-trained weights in SVD to initialize the denoising U-Net model and find that using pre-trained weights helps the model converge faster. To get better prediction quality around the robot gripper, which is important for action extraction later, we apply a re-weighting mechanism in the diffusion loss. Concretely, we use binary masks (in simulation, the object segmentations are provided; in real world, we use SAM2 [37]) of the robot gripper region and downsample it by a factor of 8 to match the resolution of the latent space while still keeping the spatial correspondence. The resulting downsampled masks provide a spatial weight map at each timestep $t^{\prime}$, denoted as $w_{g}(t^{\prime})$, which is incorporated into the joint diffusion loss mentioned in §3.3 as follows:
$\mathcal{L}=\sum_{t^{\prime}=t+1}^{t+h}\left[\left(1+\mathbbm{1}_{\{w_{g}(t^{\prime})=1\}}\right)\cdot\left(\underbrace{\mathcal{L}_{\text{diff}}^{n}(t^{\prime})+\mathcal{L}_{\text{diff}}^{m}(t^{\prime})}_{\text{RGB loss}}+\lambda\cdot\underbrace{\mathcal{L}_{\text{3D-diff}}(t^{\prime})}_{\text{pointmap loss}}\right)\right]$ (4)
where $\mathbbm{1}_{\{w_{g}(t^{\prime})=1\}}$ is an indicator function that activates if the pixel value on the spatial weight map is 1 and 0 otherwise. We add this indicator to a base weight of 1, effectively doubling the contribution of loss terms at gripper regions. This weighting encourages higher prediction accuracy in areas critical for gripper pose estimation in the policy extraction phase.
A.2 Dataset Details
To sample cameras for rendering multi-view RGBD videos, we first sample camera positions within a half-sphere shell defined by an inner radius ($r_{1}=0.7m$) and outer radius ($r_{2}=1.2m$), with the center being the origin of the world coordinate system (center of the table). We restrict the range of the camera positions within the area between $0.2m\leq x\leq 0.6m$, $-0.5m\leq y\leq 0.5m$, and $0.7\leq z\leq 1.2m$, as shown in Figure 7 (a-c). The world coordinate system is shown in Figure 7 (d).
Figure 7: Camera Sampling Visualization. We randomly sample 16 camera poses per episode using our proposed technique. (a)-(c) show example camera poses for each task, with green cameras used for training and red for evaluation. (d) shows the simulation world coordinate frame.
A.3 Training Details
The 4D generation model described in § 3 is trained separately for each task in § 4 for approximately 60 epochs using 4 NVIDIA RTX A6000 GPUs (48GB memory each). We fine-tune the full U-Net backbone of the SVD [11] model with a learning rate of $1\times 10^{-6}$, using the AdamW optimizer ($\beta_{1}=0.95$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay $=1\times 10^{-6}$) and a batch size of 4. The image and pointmap VAE encoders are frozen during diffusion model training.
At inference time, we apply the standard EulerEDMSampler [52] with 25 denoising steps. For robot policy deployment, both the generation model and the pose tracking model are run on a single NVIDIA GeForce RTX 4090 GPU.
A.4 Additional Results
Figure 8: Qualitative Multi-View Video Generation Results. We show temporal results generated by our 4D video generation model across three robot manipulation tasks. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity.
Figure 9: Qualitative Results of PlaceAppleFromBowlIntoBin task. Our method achieves the best RGB video and depth generation quality, with high multi-view consistency. Baseline results often exhibit significant cross-view inconsistencies (marked in red) or contain noticeable artifacts in the RGB or depth predictions.
Figure 10: Qualitative Results of PutSpatulaOnTable task. Our method achieves the best RGB video and depth generation quality, with high multi-view consistency. Baseline results often exhibit significant cross-view inconsistencies (marked in red) or contain noticeable artifacts in the RGB or depth predictions.
In Figure 8, we show generated RGB video sequences for the StoreCerealBoxUnderShelf, PutSpatulaOnTable, and PlaceAppleFromBowlIntoBin tasks using our proposed 4D generation model. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity. We also show baseline comparison results on the PlaceAppleFromBowlIntoBin task in Figure 9 and PutSpatulaOnTable task in Figure 10. Our method consistently achieves the best RGB video and depth generation quality, with high multi-view consistency. Baseline results often exhibit significant cross-view inconsistencies (marked in red) or contain noticeable artifacts in the RGB or depth predictions.
A.5 Broader Impact
Positive impacts. Our 4D video generation model can improve robot manipulation by providing consistent, multi-view RGB-D predictions for pose tracking, and provide better interpretability comparing to typical behavior cloning approaches. The model can be used to enhance the capabilities of general-purpose robotic systems in household, factories, etc.
Negative impacts. Like other generative models, our method could be misused for creating realistic but deceptive content (e.g., deepfakes). If used in critical settings without robust validation, errors in spatial prediction could lead to unsafe robot behaviors. Dataset biases may also limit generalization.
Mitigations. We recommend controlled model release, output watermarking, and bias-aware dataset design. In robotics, deployment should include safety checks and fallback mechanisms. Future work should explore interpretability and robustness under real-world conditions.